An investor is trying to decide between a muni paying 4 percent or an equivalent taxable corporate paying 7 percent. What is the marginal tax rate the investor must have to consider buying the municipal bond? Group of answer choices 42.85% 40% 33.85% 33.5%
Answer #1

Municipal bonds are tax free. If the after tax return of the corporate bond is greater than 4% then the investor will buy the corporate bond instead of municipal bond. Therefore, to prefer buying the municipal bond, the after tax return should be lower than 4%

For 40% tax rate, the after tax return = (1-0.40)x7 =4.2% which is greater than 4%

So the correct option has to be 42.85%

Let us verify the same: (1-0.4285) x 7= 4% so this is the cut off rate and if the rate goes higher, the investor will prefer the municipal bond. So 42.85% is the correct option.